# Data-Driven Distributed Output Synchronization of Heterogeneous Discrete-Time Multi-Agent Systems

### v1

@@ I Introduction
We use ${\rm im}(Q)$ to represent the column space of $Q$. The spectrum of a square matrix $Q$ is denoted by $\sigma(Q)$ and is the set of all its eigenvalues. The Kronecker product is denoted by $\otimes$. Given matrices $M_{i},i\in[1,p]$, the block diagonal matrix whose $i$th diagonal block is the matrix $M_{i}$ is denoted either by ${\rm diag}(M_{i}),i\in[1,p],$ or by ${\rm diag}(M_{1},M_{2},\dots,M_{p})$, while given vectors $v_{i},i\in[1,p]$, the column stacking of these vectors is denoted either by ${\rm col}(v_{i}),i\in[1,p],$ or by ${\rm col}(v_{1},v_{2},\dots,v_{p})$. The $(i,j)$th entry of a matrix $M$ is denoted by $[M]_{i,j}$.
A weighted, directed graph (digraph) is a triple ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$, where ${\mathcal{V}}=\{1,\dots,N\}=[1,N]$ is the set of nodes, ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges, and ${\mathcal{A}}\in\mathbb{R}^{N\times N}$ is the nonnegative, weighted adjacency matrix which satisfies $[{\mathcal{A}}]_{i,j}>0$ if and only if (iff) $(j,i)\in{\mathcal{E}}$. We assume that $[{\mathcal{A}}]_{i,i}=0$ for every $i\in{\mathcal{V}}$. The in-degree of the node $i$ is $d_{i}=\sum_{j=1}^{N}[{\mathcal{A}}]_{i,j}$. The in-degree matrix $\mathcal{D}$ is the diagonal matrix defined as $\mathcal{D}={\rm diag}(d_{i}),i\in[1,N]$. The Laplacian associated with ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$ is defined as $\mathcal{L}\doteq\mathcal{D}-\mathcal{A}$. A digraph is connected if there is a (directed) path from every node to every other node. Given a digraph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$, a directed spanning tree is a subgraph of ${\mathcal{G}}$ that includes all the vertices ${\mathcal{V}}$ and has a single root node from which all the other nodes can be reached, without forming any cycle.
@@ II Error-Feedback Output Synchronization: Problem Statement
[14, 15, 12] All the eigenvalues of $S$ are simple and lie on the unit circle.
@@
Consider a multi-agent system consisting of $N$ heterogeneous agents. Without loss of generality, we assume that the first $N_{l}$ agents (in the following referred to as leaders) are affected by the exosystem output $y_{0}(t)$, while the remaining $N_{f}\doteq N-N_{l}$ agents (the followers) have not. We let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$ denote the directed graph describing the interactions among the $N$ agents, and by ${\mathcal{G}}_{0}=(\{0\}\cup{\mathcal{V}},{\mathcal{E}}_{0},{\mathcal{A}}_{0})$ the extended digraph including also the node 00 corresponding to the exosystem. The set ${\mathcal{E}}_{0}$ is obtained by adding to the edges in ${\mathcal{E}}$ the edges from the exosystem node to the $N_{l}$ nodes representing the leaders. We assume that the weights of all such edges are unitary. Under the previous assumptions, the adjacency matrix ${\mathcal{A}}_{0}$ is uniquely identified from ${\mathcal{A}}$, and the Laplacian matrix associated with it can be expressed in partitioned form as follows:
$\mathcal{L}_{0}=\begin{bmatrix}0&\mathbb{0}_{N_{l}}^{\top}&\mathbb{0}_{N_{f}}^{\top}\\ \mathbb{1}_{N_{l}}&\mathcal{L}_{ll}&\mathcal{L}_{lf}\\ \mathbb{0}_{N_{f}}&\mathcal{L}_{fl}&\mathcal{L}_{ff}\\ \end{bmatrix}$ (2)
@@
[3, 14, 15, 12] The digraph ${\mathcal{G}_{0}}$ contains a directed spanning tree with root node 00.
@@
where $[{\mathcal{A}}]_{i,j}$ is the $(i,j)$th entry of the adjacency matrix ${\mathcal{A}}$, while $H\in{\mathbb{R}}^{n_{0}\times p}$ is a matrix parameter to be designed.
@@ III-A Leader Nodes Dynamics
$e_{l}(t)=$ $(C_{l}+K_{l}D_{l})\varepsilon_{l}(t)+(C\Pi_{l}+D_{l}\Gamma_{l})\delta_{l}(t)$ (16c)
@@
Note that the leader dynamics is completely decoupled from the follower dynamics. By resorting to a change of basis we can replace the state vector $[x_{l}(t)^{\top}\ z_{l}(t)^{\top}\ ({\mathbb{1}}_{N_{l}}\otimes x_{0}(t))^{\top}]^{\top}$ with $[\delta_{l}(t)^{\top}\ \varepsilon_{l}(t)^{\top}\ ({\mathbb{1}}_{N_{l}}\otimes x_{0}(t))^{\top}]^{\top}$. Consequently, under Assumptions 1, 2, and 3, we can ensure that the internal dynamics of the leaders are asymptotically stable when $x_{0}(t)$ is identically zero (see point 1) of Problem 1) and the error $e_{l}(t)\to\mathbb{0}$ as $t\to\infty$ (see point 2) of Problem 1) iff
@@
$e_{l}(t)=$ $(C_{l}+K_{l}D_{l})\varepsilon_{l}(t)+(C\Pi_{l}+D_{l}\Gamma_{l})\delta_{l}(t).$ (18c)
@@ III-B Follower Nodes Dynamics
Note that the follower dynamics is decoupled, in turn, from the leader dynamics. By resorting to a change of basis we replace the state vector $[x_{f}(t)^{\top}\ z_{f}(t)^{\top}\ ({\mathbb{1}}_{N_{f}}\otimes x_{0}(t))^{\top}]^{\top}$ with $[\delta_{f}(t)^{\top}\ \varepsilon_{f}(t)^{\top}\ ({\mathbb{1}}_{N_{f}}\otimes x_{0}(t))^{\top}]^{\top}$. So, under the proposed assumptions, we can ensure that the internal dynamics of the followers are asymptotically stable (see point 1) of Problem 1) and $e_{f}(t)\to{\mathbb{0}}$ as $t\to\infty$ (see point 2) of Problem 1) iff
@@ III-C Model-Based Solution
We observe, now, that there exist matrices $K_{i},i\in[1,N]$, such that conditions c1) and c4) hold if and only if each pair $(A_{i},B_{i})$ is stabilizable, which is exactly condition $i)$. Finally, conditions c3) and c6) are equivalent to $ii)$ and $iii)$.
@@ IV Data-Driven Approach
In order to solve Problem 1 in a data-driven set-up, we need to understand how conditions $i)$, $ii)$ and $iii)$ in Theorem 1 translate in terms of data matrices. Note that while condition $ii)$ pertains only leaders and condition $iii)$ only followers, condition $i)$ pertains both. However, due to the different state-space models generating the collected data, the conditions we will derive for leaders and followers will be different.
@@ IV-A Leader Nodes Dynamics
1)2)1$\Leftrightarrow$ 2)1 ) ⇔ 2 )  From the previous analysis, it follows that condition 1) holds if and only if there exists $Q_{i}$ such that
@@
2)3)2$\Leftrightarrow$ 3)2 ) ⇔ 3 )  Follows from the PBH reachability test [19].
We now address condition $ii)$ in Theorem 1 from a data-driven perspective.
@@
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ) If we denote by $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})$ the “real” sextuple that generated the data, then clearly $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})\in\Sigma_{i}^{l}$ and the sextuples $(\hat{A}_{i},\hat{B}_{i},\hat{E}_{i},\hat{C}_{i},\hat{D}_{i},\hat{F}_{i})$ compatible with the data are those and only those that satisfy
@@
${\rm ker_{L}}\left(\begin{bmatrix}X_{i}^{p}\\ U_{i}^{p}\\ Y_{0}^{p}\end{bmatrix}\right)\subseteq{\rm ker_{L}}\left(\begin{bmatrix}\Pi_{i}\\ \Gamma_{i}\\ R\end{bmatrix}\right)$
@@
2)1)2$\Rightarrow$ 1)2 ) ⇒ 1 ) Suppose that there exists a matrix $M_{i}$ such that (32) hold. Set $\Pi_{i}\doteq X_{i}^{p}M_{i},$ and $\Gamma_{i}\doteq U_{i}^{p}M_{i}.$ For every sextuple $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})\in\Sigma_{i}^{l}$, condition (32a), making use of (32c), leads to
@@ IV-B Follower Nodes Dynamics
Note that corresponding to every $(X_{i}^{p})^{\#}$ satisfying $ii)$, we obtain the stabilizing feedback gain $K_{i}=U_{i}^{p}(X_{i}^{p})^{\#}$ (see [22]). The following result is the analogous, for followers, of Proposition 5 and hence its proof is omitted.
@@ IV-C Data-driven solution
Consider an exosystem and a group of 5 agents (2 leaders and 3 followers) connected through the binary (i.e. $[\mathcal{A}]_{ij}$ is either 00 or $1$) digraph $\mathcal{G}_{0}$ depicted in Figure 1 and with dynamics described by the following matrices:
@@ V Conclusions
a) Consider the condensed graph ${\mathcal{G}}_{c}$, obtained from ${\mathcal{G}}_{0}$ by merging all leader nodes and the node 00 representing the exosystem. There is an edge from this new node to any of the follower nodes if and only if there was an edge from one of the leaders to that follower in ${\mathcal{G}}_{0}$. Conversely, there is an edge from any follower to the new node if and only if there was an edge from that follower to one of the leaders in ${\mathcal{G}}_{0}$. It is easy to see that the Laplacian associated with this new digraph is related to ${\mathcal{L}}_{0}$ in (2) as follows:
${\mathcal{L}}_{c}=\begin{bmatrix}N_{l}+{\mathbb{1}}_{N_{l}}^{\top}{\mathcal{L}}_{ll}{\mathbb{1}}_{N_{l}}&{\mathbb{1}}_{N_{l}}^{\top}{\mathcal{L}}_{lf}\cr{\mathcal{L}}_{fl}{\mathbb{1}}_{N_{l}}&{\mathcal{L}}_{ff}.\end{bmatrix}$
@@
b) By Gershgorin’s Circle theorem [11], we can say that
$\sigma$ $((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})$
$\subset\bigcup_{j\in[N_{l}+1,N]}\left\{z\in\mathbb{C}\ :\ \left|z-\frac{d_{j}}{1+d_{j}}\right|\leq\sum_{k=N_{l}+1}^{N}\frac{[\mathcal{A}]_{jk}}{1+d_{j}}\right\},$
but, by Assumption 3, $\sum_{k=N_{l}+1}^{N}[\mathcal{A}]_{jk}\leq d_{j}$, and hence for all $j\in[N_{l}+1,N]$ and every $\lambda\in\sigma((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})$:
@@
that, together with a), implies b).
@@
Set $\sigma((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})=\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{f}}\}$. Under Assumptions 1, 2 and 3, there exists a matrix $H$ such that $S-\lambda_{i}HR$ is Schur stable for all $i\in[1,N_{f}]$.
@@
First, we want to prove that if $(R,S)$ is observable, then there exists a vector $v\neq 0$ such that $(v^{\top}R,S)$ is observable. Let $T$ be a nonsingular matrix s.t. $T^{-1}ST={\rm diag}(\mu_{1},\dots,\mu_{n_{0}})$, with $|\mu_{j}|=1$, $\mu_{k}\neq\mu_{j}$ if $k\neq j$. It follows that $(R,S)$ is observable if and only if $(RT,T^{-1}ST)$ is observable, and by the PBH observability criterion this is the case if and only if $RT$ does not have null columns. On the other hand, $(v^{\top}R,S)$ is observable if and only if $(v^{\top}RT,T^{-1}ST)$ is observable, which happens if and only if $v^{\top}RT$ does not have null entries.
Let $w_{i}$ denote the $i$th column of $RT$, then $v^{\top}w_{i}=0$ if and only if $v\in\left({\rm im}(w_{i})\right)^{\perp}$. Since $\bigcup_{i=1}^{n_{0}}\left({\rm im}(w_{i})\right)^{\perp}\subsetneqq\mathbb{R}^{p}$, it follows that exists $v\in\mathbb{R}^{p}\setminus\left(\bigcup_{i=1}^{n_{0}}\left({\rm im}(w_{i})\right)^{\perp}\right)$, namely exists $v\neq\mathbb{0}$ such that $v^{\top}RT$ does not have null entries, and thus $(v^{\top}R,S)$ is observable.
Now it remains to show that if $(v^{\top}R,S)$ is observable then there exists $q\in{\mathbb{R}}^{n_{0}}$ such that $S-\lambda_{i}qv^{\top}R$ is Schur for all $i\in[1,N_{f}]$, but this is the dual result of [23, Theorem 3.2]. Note that we can apply such result because condition (17) in [23] holds, due to Lemma 11. So, the result holds for $H=qv^{\top}.$
@@
Let $A\in{\mathbb{R}}^{n\times r}$ and $B\in{\mathbb{R}}^{k\times r}$, and assume that
$\Psi\doteq\begin{bmatrix}A\cr B\end{bmatrix}\in{\mathbb{R}}^{(n+k)\times r}$
is of full row rank. A matrix $C\in{\mathbb{R}}^{r\times n}$ satisfies
$\Psi C=\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$ (35)
if and only if
$C=\left(\Psi^{\dagger}+(I_{r}-\Psi^{\dagger}\Psi)Q\right)\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$ (36)
for some matrix $Q\in{\mathbb{R}}^{r\times(n+k)}$.

### v2

@@ I Introduction
We use ${\rm im}(Q)$ to represent the column space of $Q$. The spectrum of a square matrix $Q$ is denoted by $\sigma(Q)$ and is the set of all its eigenvalues. The Kronecker product is denoted by $\otimes$. Given matrices $M_{i},i\in[1,p]$, the block diagonal matrix whose $i$th diagonal block is the matrix $M_{i}$ is denoted either by ${\rm diag}(M_{i}),i\in[1,p],$ or by ${\rm diag}(M_{1},M_{2},\dots,M_{p})$, while given vectors $v_{i},i\in[1,p]$, the column stacking of these vectors is denoted either by ${\rm col}(v_{i}),i\in[1,p],$ or by ${\rm col}(v_{1},v_{2},\dots,v_{p})$. The $(i,j)$th entry of a matrix $M$ is denoted by $M]_{i,j}$.
A weighted, directed graph (digraph) is a triple ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$, where ${\mathcal{V}}=\{1,\dots,N\}=[1,N]$ is the set of nodes, ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges, and ${\mathcal{A}}\in\mathbb{R}^{N\times N}$ is the nonnegative, weighted adjacency matrix which satisfies ${\mathcal{A}}]_{i,j}>0$ if and only if (iff) $(j,i)\in{\mathcal{E}}$. We assume that ${\mathcal{A}}]_{i,i}=0$ for every $i\in{\mathcal{V}}$. The in-degree of the node $i$ is $d_{i}=\sum_{j=1}^{N}[{\mathcal{A}}]_{i,j}$. The in-degree matrix $\mathcal{D}$ is the diagonal matrix defined as $\mathcal{D}={\rm diag}(d_{i}),i\in[1,N]$. The Laplacian associated with ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$ is defined as $\mathcal{L}\doteq\mathcal{D}-\mathcal{A}$. A digraph is connected if there is a (directed) path from every node to every other node. Given a digraph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$, a directed spanning tree is a subgraph of $\mathcal{G}$} that includes all the vertices $\mathcal{V}$} and has a single root node from which all the other nodes can be reached, without forming any cycle.
@@ II Error-Feedback Output Synchronization: Problem Statement
[14, 15] All the eigenvalues of $S$ are simple and lie on the unit circle.
@@
Consider a multi-agent system consisting of $N$ heterogeneous agents. Without loss of generality, we assume that the first $N_{l}$ agents (in the following referred to as leaders) are affected by the exosystem output $y_{0}(t)$, while the remaining $N_{f}\doteq N-N_{l}$ agents (the followers) have not. We let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{A}})$ denote the directed graph describing the interactions among the $N$ agents, and by ${\mathcal{G}}_{0}=(\{0\}\cup{\mathcal{V}},{\mathcal{E}}_{0},{\mathcal{A}}_{0})$ the extended digraph including also the node 0 corresponding to the exosystem. The set ${\mathcal{E}}_{0}$ is obtained by adding to the edges in $\mathcal{E}$} the edges from the exosystem node to the $N_{l}$ nodes representing the leaders. We assume that the weights of all such edges are unitary. Under the previous assumptions, the adjacency matrix ${\mathcal{A}}_{0}$ is uniquely identified from $\mathcal{A}$}, and the Laplacian matrix associated with it can be expressed in partitioned form as follows:
$\mathcal{L}_{0}=\begin{bmatrix}0&\mathbb{0}_{N_{l}}^{\top}&\mathbb{0}_{N_{f}}^{\top}\\ {-}\mathbb{1}_{N_{l}}&\mathcal{L}_{ll}&\mathcal{L}_{lf}\\ \mathbb{0}_{N_{f}}&\mathcal{L}_{fl}&\mathcal{L}_{ff}\\ \end{bmatrix}$ (2)
@@
[3, 14, 15, 12] The digraph $\mathcal{G}_{0}$} contains a directed spanning tree with root node 0.
@@
where ${\mathcal{A}}]_{i,j}$ is the $(i,j)$th entry of the adjacency matrix $\mathcal{A}$}, while $H\in{\mathbb{R}}^{n_{0}\times p}$ is a matrix parameter to be designed.
@@ III-A Leader Nodes Dynamics
$e_{l}(t)=$ $(C_{l}+{D_{l}K_{l}})\varepsilon_{l}(t)+(C\Pi_{l}+D_{l}\Gamma_{l})\delta_{l}(t)$ (16c)
@@
Note that the leader dynamics is completely decoupled from the follower dynamics. By resorting to a change of basis we can replace the state vector $x_{l}(t)^{\top}\ z_{l}(t)^{\top}\ ({\mathbb{1}}_{N_{l}}\otimes x_{0}(t))^{\top}]^{\top}$ with $\delta_{l}(t)^{\top}\ \varepsilon_{l}(t)^{\top}\ ({\mathbb{1}}_{N_{l}}\otimes x_{0}(t))^{\top}]^{\top}$. Consequently, under Assumptions 1, 2, and 3, we can ensure that the internal dynamics of the leaders are asymptotically stable when $x_{0}(t)$ is identically zero (see point 1) of Problem 1) and the error $e_{l}(t)\to\mathbb{0}$ as $t\to\infty$ (see point 2) of Problem 1) iff
@@
$e_{l}(t)=$ $(C_{l}+{D_{l}K_{l}})\varepsilon_{l}(t)+(C\Pi_{l}+D_{l}\Gamma_{l})\delta_{l}(t).$ (18c)
@@ III-B Follower Nodes Dynamics
Note that the follower dynamics is decoupled, in turn, from the leader dynamics. By resorting to a change of basis we replace the state vector $x_{f}(t)^{\top}\ z_{f}(t)^{\top}\ ({\mathbb{1}}_{N_{f}}\otimes x_{0}(t))^{\top}]^{\top}$ with $\delta_{f}(t)^{\top}\ \varepsilon_{f}(t)^{\top}\ ({\mathbb{1}}_{N_{f}}\otimes x_{0}(t))^{\top}]^{\top}$. So, under the proposed assumptions, we can ensure that the internal dynamics of the followers are asymptotically stable (see point 1) of Problem 1) and $e_{f}(t)\to{\mathbb{0}}$ as $t\to\infty$ (see point 2) of Problem 1) iff
@@ III-C Model-Based Solution
We observe, now, that there exist matrices $K_{i},i\in[1,N]$, such that conditions c1) and c4) hold if and only if each pair $(A_{i},B_{i})$ is stabilizable, which is exactly condition $i)$. Finally, conditions c3) and c6) are equivalent to ii)ii) and $iii)$.
@@ IV Data-Driven Approach
In order to solve Problem 1 in a data-driven set-up, we need to understand how conditions $i)$, ii)ii) and $iii)$ in Theorem 1 translate in terms of data matrices. Note that while condition ii)ii) pertains only leaders and condition $iii)$ only followers, condition $i)$ pertains both. However, due to the different state-space models generating the collected data, the conditions we will derive for leaders and followers will be different.
@@ IV-A Leader Nodes Dynamics
$1)\Leftrightarrow 2)$  From the previous analysis, it follows that condition 1) holds if and only if there exists $Q_{i}$ such that
@@
$2)\Leftrightarrow 3)$  Follows from the PBH reachability test [19].
We now address condition ii)ii) in Theorem 1 from a data-driven perspective.
@@
$1)\Rightarrow 2)$ If we denote by $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})$ the “real” sextuple that generated the data, then clearly $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})\in\Sigma_{i}^{l}$ and the sextuples $(\hat{A}_{i},\hat{B}_{i},\hat{E}_{i},\hat{C}_{i},\hat{D}_{i},\hat{F}_{i})$ compatible with the data are those and only those that satisfy
@@
[…]
@@
$2)\Rightarrow 1)$ Suppose that there exists a matrix $M_{i}$ such that (32) hold. Set $\Pi_{i}\doteq X_{i}^{p}M_{i},$ and $\Gamma_{i}\doteq U_{i}^{p}M_{i}.$ For every sextuple $(A_{i},B_{i},E_{i},C_{i},D_{i},F_{i})\in\Sigma_{i}^{l}$, condition (32a), making use of (32c), leads to
@@ IV-B Follower Nodes Dynamics
Note that corresponding to every $(X_{i}^{p})^{\#}$ satisfying ii)ii), we obtain the stabilizing feedback gain $K_{i}=U_{i}^{p}(X_{i}^{p})^{\#}$ (see [22]). The following result is the analogous, for followers, of Proposition 5 and hence its proof is omitted.
@@ IV-C Data-driven solution
Consider an exosystem and a group of 5 agents (2 leaders and 3 followers) connected through the binary (i.e. $\mathcal{A}]_{ij}$ is either 0 or $1$) digraph $\mathcal{G}_{0}$ depicted in Figure 1 and with dynamics described by the following matrices:
@@ Appendix A Appendix
a) Consider the condensed graph ${\mathcal{G}}_{c}$, obtained from ${\mathcal{G}}_{0}$ by merging all leader nodes with the node 0 representing the exosystem. There is an edge from this new node to any of the follower nodes iff there was an edge from one of the leaders to that follower in ${\mathcal{G}}_{0}$. Conversely, there is an edge from any follower to the new node iff there was an edge from that follower to one of the leaders in ${\mathcal{G}}_{0}$. It is easy to see that the Laplacian associated with this new digraph is related to ${\mathcal{L}}_{0}$ in (2) as follows:
${\mathcal{L}}_{c}=\begin{bmatrix}-N_{\ell}+{\mathbb{1}}_{N_{\ell}}^{\top}{\mathcal{L}}_{\ell\ell}{\mathbb{1}}_{N_{\ell}}&{\mathbb{1}}_{N_{\ell}}^{\top}{\mathcal{L}}_{\ell f}\cr{\mathcal{L}}_{f\ell}{\mathbb{1}}_{N_{\ell}}&{\mathcal{L}}_{ff}.\end{bmatrix}$
@@
b) By Gershgorin’s Circle theorem [11], we can say that $\sigma((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})$ is included in $\bigcup_{j\in[N_{\ell}+1,N]}\left\{z\in\mathbb{C}\ :\ \left|z-\frac{d_{j}}{1+d_{j}}\right|\leq\sum_{k=N_{\ell}+1}^{N}\frac{[\mathcal{A}]_{jk}}{1+d_{j}}\right\},$ but $\sum_{k=N_{\ell}+1}^{N}[\mathcal{A}]_{jk}\leq d_{j}$, and hence for all $j\in[N_{\ell}+1,N]$ and every $\lambda\in\sigma((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})$:
@@
that, together with a), implies b). ∎
@@
Set $\sigma((I_{N_{f}}+\mathcal{D}_{f})^{-1}\mathcal{L}_{ff})=\{\lambda_{1},\dots,\lambda_{N_{f}}\}$. Under Assumptions 1, 2 and 3, there exists a matrix $H$ such that $S-\lambda_{k}HR$ is Schur stable for all $k\in[1,N_{f}]$.
@@
First, we want to prove that if $(R,S)$ is observable, then there exists a vector $v\neq\mathbb{0}$ such that $(v^{\top}R,S)$ is observable. Let $W$ be a nonsingular matrix s.t. $W^{-1}SW={\rm diag}(\mu_{1},\dots,\mu_{n_{0}})$, with $|\mu_{j}|=1$, $\mu_{i}\neq\mu_{j}$ if $i\neq j$. It follows that $(R,S)$ is observable iff $(RW,W^{-1}SW)$ is observable, and by the PBH observability criterion this is the case iff $RW$ does not have null columns. On the other hand, $(v^{\top}R,S)$ is observable iff $(v^{\top}RW,W^{-1}SW)$ is observable, which happens iff $v^{\top}RW$ does not have null entries.
Let $\bar{w}_{i}$ denote the $i$th column of $RW$, then $v^{\top}\bar{w}_{i}=0$ iff $v\in\left({\rm im}(\bar{w}_{i})\right)^{\perp}$. Since $\bigcup_{i=1}^{n_{0}}\left({\rm im}(\bar{w}_{i})\right)^{\perp}\subsetneqq\mathbb{R}^{p}$, it follows that exists $v\in\mathbb{R}^{p}\setminus\left(\bigcup_{i=1}^{n_{0}}\left({\rm im}(\bar{w}_{i})\right)^{\perp}\right)$, namely exists $v\neq\mathbb{0}$ such that $v^{\top}RW$ does not have null entries, and thus $(v^{\top}R,S)$ is observable.
Now it remains to show that if $(v^{\top}R,S)$ is observable then there exists $q\in{\mathbb{R}}^{n_{0}}$ such that $S-\lambda_{k}qv^{\top}R$ is Schur for all $k\in[1,N_{f}]$, but this is the dual result of [23, Theorem 3.2]. Note that we can apply such result because condition (17) in [23] holds, due to Lemma 11. So, the result holds for $H=qv^{\top}.$
@@
Let $A\in{\mathbb{R}}^{n\times r}$ and $B\in{\mathbb{R}}^{k\times r}$, with ${\rm rank}(A)=n$ and ${\rm rank}(B)=\bar{k}$, and suppose that
$\bar{\Psi}\doteq\begin{bmatrix}A\cr B\end{bmatrix}\in{\mathbb{R}}^{(n+k)\times r}$
has ${\rm rank}(\bar{\Psi})=n+{\bar{k}}.$ A matrix $C\in{\mathbb{R}}^{r\times n}$ satisfies
$\bar{\Psi}C=\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$ (35)
if and only if there exists $Q\in{\mathbb{R}}^{r\times(n+k)}$ such that
$C=\left(\bar{\Psi}^{\dagger}+(I_{r}-\bar{\Psi}^{\dagger}\bar{\Psi})Q\right)\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$. (36)